OpenStudy (anonymous):

A wire carries a 1.7A current for 10s . How many electrons move past a point in the wire? Ne=? I have tried this, I feel my math is correct, any other ideas.

OpenStudy (doc.brown):

\[A = \frac{6.241\times 10^{18}electrons}{second}\]\[1.7\times\frac{6.241\times 10^{18}e}{s}\times 10s\]\[=1.1\times 10^{20}e\]

OpenStudy (doc.brown):

@ GopZ Q is electric charge.

OpenStudy (anonymous):

@doc, yeah m aware of that, q=I.t ryt?

OpenStudy (anonymous):

Q=I.t =10*1,7 =17 Ne=17/1,6*10^-19 =1,1*10^20 electrons @doc, my mistake, thanks! #wink

OpenStudy (doc.brown):

Cool